OpenStudy (anonymous):

Find the slope intercept equation of the line passing through the points (3/4, 3/2) and (-4/3,7/4)

OpenStudy (campbell_st):

find the gradient rise/run or m = (7/4 - 3/2)/(-4/3 - 3/4) then use point gradient formula \[y - y _{1} = m(x - x _{2})\] with either point

OpenStudy (anonymous):

use this formula for intercept equation of the line.  such that the vertices given indicate (a,b)